# The seeding method: A test case for classical nucleation theory in small systems

### v1

@@ 1 Introduction
Nucleation is a rare and thermally activated process. In molecular dynamics (MD) simulations, only small critical clusters with low nucleation barrier — in the regime of high supersaturation — are accessible with brute-force approaches [1, 2, 3, 4, 5, 6, 7], even when the simulation system is large [8, 9, 10]. While nucleation rates can be readily measured from brute-force simulations [11, 1], identifying critical clusters remains highly challenging [12]. A common strategy to estimate the critical cluster size is to apply the nucleation theorem [13] from measured nucleation rate [8]. However, in the high-supersaturation regime, classical nucleation assumptions become questionable [7]. To overcome these limitations, a variety of rare-event simulation techniques have been developed to promote the formation of critical cluster [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. These methods are well suited for investigating nucleation with larger nucleation barriers and at lower supersaturation [28]. In parallel, another approach has emerged: the seeding technique [29, 30, 31, 32]. In seeded MD simulations, the system is initialized with a pre-formed nucleus of the new phase. At constant pressure and temperature (NPT ensemble), the inserted seed may either grow (if post-critical) or redissolve (if pre-critical). At the exact critical size, the cluster corresponds to an unstable equilibrium, with a $50\%$ probability of growth or dissolution. Once the critical cluster is identified, classical nucleation theory (CNT) [13, 33, 34, 35, 36, 37, 38], on which the approach relies, can then be applied to estimate the nucleation rate [29]. This approach is conceptually simple and intuitive, but in practice requires many independent simulations to pinpoint the condition at which the seed is critical. Moreover, because the seeds typically grow or redissolve rapidly, NPT seeding can be computationally intensive when aiming for robust statistics.
@@ 2 Simulation details
Simulations were performed using the open-source package LAMMPS [52, 53]. For the Lennard-Jones interactions, we used a large cutoff of $6.78\sigma$ with $\sigma$ the length parameter [54, 55]. Lennard-Jones units are used in the following. Energies are scaled by $\epsilon$, which is the energy depth of the untruncated Lennard-Jones potential. Lengths are scaled by $\sigma$, temperatures by $\epsilon/k_{\mathrm{B}}$, densities by $\sigma^{-3}$, and time by $\tau=\sqrt{m\sigma^{2}/\epsilon}$ with $m$ the mass of the particle. The timestep employed is $0.005\tau$. Cubic periodic boundary conditions are used. For temperature control, we employed the Nose-Hoover thermostat implemented in LAMMPS with damping parameter taken as $0.5\tau$. We use the following procedure for performing seeded simulations. The first step is to prepare the seed. The seed is obtained from a NVT run in a box of arbitrary size to equilibrate the single phase liquid state at the density of the phase diagram $\bar{\rho}_{l}$, which sets the number of particles in the chosen volume. Then a sphere of a given size $R$ is extracted from the liquid phase and inserted in a cubic box of size $L$. This sets the initial number of liquid particles, $N_{l}$, in the seeded simulation:
@@ 3 Classical nucleation theory
In practice, the critical densities of both the stable and unstable clusters are very close to that of the phase diagram, $\bar{\rho}_{l}$. If we further assume that the vapor phase is ideal and the liquid phase incompressible, one gets the usual relation for both critical sizes,
@@
in reduced units, with $S^{*}=p_{v}^{*}/\bar{p}_{v}$ the ratio between the vapor pressure at equilibrium and that of the saturated vapor. Since the vapor is assumed to be ideal gas, $S^{*}=\rho_{v}^{*}/\bar{\rho}_{v}$, where $\bar{\rho}_{v}$ is the saturated vapor density of the phase diagram. Together with Eq. (6) and $\rho_{l}^{*}=\bar{\rho}_{l}$, Eq. (7) constitutes the approximate problem derived by Reguera et al. [45] to find both the unstable and stable clusters in confined systems. However, if quantitative predictions are intended, a more rigorous thermodynamic description of the chemical potential should be required, as we shall see.
@@ 4 Results
We investigate the condensation of Lennard-Jones particles over the temperature range from $0.7$ to $1.1$. The two solutions of Eqs. (4), (5), and (6) are obtained using chemical potential computed from the EOS developed by Johnson, Zollweg, and Gubbins (JZG) [49]. The EOS are combined with a mean-field correction for the cut potential [49], using the same cutoff ($6.78$) as in the simulations. Within this temperature range, the phase diagram predicted by the JZG EOS is found to be in excellent agreement with that obtained from MD simulations of the two-phase system at equilibrium (not shown). The surface tension of the planar vapor/liquid interface at equilibrium is classically calculated from MD simulations using the Kirkwood pressure tensor method [57]. The results follow the expected power law [55], $\gamma=\gamma_{0}\left(1-T/T_{c}\right)^{a}$, with $\gamma_{0}=2.838$, $a=1.252$, and $T_{c}=1.305$ the critical temperature.
@@
The smaller droplet (dashed lines) corresponds to the unstable equilibrium, while the larger droplet (solid lines) represents the stable cluster. As expected [45, 47, 48], the two branches merge for the density at which the superstabilization effect operates, i.e., no cluster forms since the vapor phase at this given box density becomes stable. For supersaturation between the binodal limit (vertical black dotted line) and this box density, nucleation is therefore impeded [45]. Predictions obtained from the approximate relations, Eq. (6) with $\rho_{l}^{*}=\bar{\rho}_{l}$ and Eq. (7), are also shown: stable and unstable branches are represented by dash-dotted and dotted lines, respectively. They agree very well with the JZG EOS results, particularly for the upper branches corresponding to the stable clusters.
@@
In order to measure the cluster properties, we use Ovito [58] and first segment the liquid droplet by clustering analysis. A clustering distance, $d$, is defined, and particles forming a connected set — each within a distance $d$ of at least one other — are identified; the number of such particles are denoted $n^{*}$. The clustering analysis is not very sensitive to $d$ within the conventional range [1.2, 1.5][1.2,\,1.5] for Lennard-Jones droplets [8]; in our analysis we adopt $d=1.25$. Several methods available in Ovito’s MD standard analysis library can then be used to determine the cluster radius: (i) fitting the radial density profile, or its integral, in simulation to a modified sigmoidal function [9, 59], (ii) the convex hull method [58], or (iii) the radius of gyration [58]. The radius may also be estimated from the droplet volume, $R^{*}=(3n^{*}/(4\mathrm{\pi}\bar{\rho}_{l}))^{1/3}$, assuming the droplet is homogeneous with bulk liquid density $\bar{\rho}_{l}$ from the phase diagram. In our case, all methods yield very similar results, since the clusters are well defined, nearly spherical, and relatively large. For sake of clarity, only the results obtained from the droplet volume estimation are shown. The corresponding results are also shown in Fig. 1(a) as light-colored circles with error bars. The error bars, representing standard deviations in $R^{*}$, are too small to be visible. With three configurations ($L=30$, $50$, and $70$), stable clusters with radius ranging from $5$ to $16$ are obtained. Remarkably, the seeding results are reproduced very well by CNT predictions using either the JZG EOS or the ideal gas approximation. Fig. 1(b)-(e) present seeded simulation results at higher temperatures: $T=0.8$, $T=0.9$, $T=1$, and $T=1.1$. Agreement with CNT predictions using the JZG EOS remains excellent across this range. In contrast, the ideal gas approximation becomes unreliable at high temperature. Nevertheless, it provides sufficiently accurate estimates to serve as practical initial conditions for seeded simulations, owing to the simplicity of the corresponding CNT calculations.
@@ 5 Summary
The seeding technique has been applied to investigate the condensation of Lennard-Jones particles and, at the same time, to test both CNT and the accuracy of widely used thermodynamic models for Lennard-Jones fluid. The properties of the critical clusters predicted by CNT are found to agree well with seeded simulation results when JZG EOS is employed. The KN EOS also performs relatively well. In contrast, caution is required when using DFT or the ideal gas approximation, as both exhibit significant inaccuracies at high temperatures ($T\geq 0.9$). Nevertheless, the ideal gas approximation in confined systems, as expressed in Eqs. (6) and (7), remains of very practical value. It provides a simple and effective framework for initializing and guiding NVT seeded simulations, especially when no thermodynamic model for chemical potential is available.

### v2

@@ 1 Introduction
Nucleation is a rare and thermally activated process. In molecular dynamics (MD) simulations, only small critical clusters with low nucleation barrier — in the regime of high supersaturation — are accessible with brute-force approaches [1, 2, 3, 4, 5, 6, 7], even when the simulation system is large [8, 9, 10]. While nucleation rates can be readily measured from brute-force simulations [11, 1], identifying critical clusters remains highly challenging [12]. A common strategy to estimate the critical cluster size is to apply the nucleation theorem [13] from measured nucleation rate [8]. However, in the high-supersaturation regime, classical nucleation assumptions become questionable [7]. To overcome these limitations, a variety of rare-event simulation techniques have been developed to promote the formation of critical cluster [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In parallel, another approach has emerged: the seeding technique [31, 32, 33, 34]. The technique was employed to investigate cavitation and condensation, as well as crystal nucleation [31, 35]. In seeded MD simulations, the system is initialized with a pre-formed nucleus of the new phase. At constant pressure and temperature (NPT ensemble), the inserted seed may either grow (if post-critical) or redissolve (if pre-critical). At the exact critical size, the cluster corresponds to an unstable equilibrium, with a $50\%$ probability of growth or dissolution. Once the critical cluster is identified, classical nucleation theory (CNT) [13, 36, 37, 38, 39, 40, 41], on which the approach relies, can then be applied to estimate the nucleation rate [31]. This approach is conceptually simple and intuitive, but in practice requires many independent simulations to pinpoint the condition at which the seed is critical. Moreover, because the seeds typically grow or redissolve rapidly, NPT seeding can be computationally intensive when aiming for robust statistics.
@@ 2 Simulation details
Simulations were performed using the open-source package LAMMPS [55, 56]. For the Lennard-Jones interactions, we used a large cutoff of $6.78\sigma$ with $\sigma$ the length parameter [57, 58]. Lennard-Jones units are used in the following. Energies are scaled by $\epsilon$, which is the energy depth of the untruncated Lennard-Jones potential. Lengths are scaled by $\sigma$, temperatures by $\epsilon/k_{\mathrm{B}}$, densities by $\sigma^{-3}$, and time by $\tau=\sqrt{m\sigma^{2}/\epsilon}$ with $m$ the mass of the particle. The timestep employed is $0.005\tau$. Cubic periodic boundary conditions are used. For temperature control, we employed the Nose-Hoover thermostat [59, 55] implemented in LAMMPS. We use the following procedure for performing seeded simulations. The first step is to prepare the seed. The seed is obtained from a NVT run in a box of arbitrary size to equilibrate the single phase liquid state at the density of the phase diagram $\bar{\rho}_{l}$, which sets the number of particles in the chosen volume. Then a sphere of a given size $R$ is extracted from the liquid phase and inserted in a cubic box of size $L$. This sets the initial number of liquid particles, $N_{l}$, in the seeded simulation:
@@ 3 Classical nucleation theory
Figure 1: Free energy associated with the formation of a spherical droplet of radius $R$ in a box of size $L$ with density $\rho=0.025$ at $T=0.8$.
In practice, the critical densities of both the stable and unstable clusters are very close to that of the phase diagram, $\bar{\rho}_{l}$. If we further assume that the vapor phase is ideal and the liquid phase incompressible, one gets in reduced units the following relation
$\frac{2\gamma}{R^{*}}=\bar{\rho}_{l}T\ln{\frac{p_{v}^{*}}{\bar{p}_{v}}}-\left(p_{v}-\bar{p}_{v}\right),$ (7)
where we have used $\mu_{l}=\bar{\mu}_{v}+(p_{l}-\bar{p}_{v})/\bar{\rho}_{l}$ with $\bar{p}_{v}$ the pressure of the saturated vapor and $\mu_{v}=\bar{\mu}_{v}+T\ln{(p_{v}/\bar{p}_{v})}$ together with Eqs.4 and 5. The last term in Eq.7 is usually very small and is commonly discarded and the well-known result is recovered for the critical size
@@
with $S^{*}=p_{v}^{*}/\bar{p}_{v}$ the ratio between the vapor pressure at equilibrium and that of the saturated vapor. Since the vapor is assumed to be ideal gas, $S^{*}=\rho_{v}^{*}/\bar{\rho}_{v}$, where $\bar{\rho}_{v}$ is the saturated vapor density of the phase diagram. Together with Eq. (6) and $\rho_{l}^{*}=\bar{\rho}_{l}$, Eq. (8) constitutes the approximate problem derived by Reguera et al. [48] to find both the unstable and stable clusters in confined systems. However, if quantitative predictions are intended, a more rigorous thermodynamic description of the chemical potential should be required, as we shall see.
Aside from the coexistence conditions, it is always useful to examine the free energy landscape for nucleation. The free energy difference of the confined system, $\Delta F=F(R)-F(0)$, associated with the formation of a spherical droplet of radius $R$ in a box of size $L$ with density $\rho$, can be expressed as
$\Delta F=-\bar{\rho}_{l}V_{l}T\ln{\frac{\rho_{v}}{\bar{\rho}_{v}}}+A_{l}\gamma+L^{3}T\left(\rho-\Phi\bar{\rho}_{v}+\rho_{v}\left(\Phi-1\right)+\rho\ln{\frac{\rho_{v}}{\rho}}\right).$ (9)
Again we assumed that the droplet is incompressible and the vapor is ideal. $A_{l}$ refers to the droplet surface. This free energy landscape is shown in Fig. 1 for various box size with global density $\rho=0.025$ at temperature $T=0.8$. The value of $\gamma$ is taken from Ref. [58]. For an infinite system, the classical result is recovered: the free energy exhibits a single maximum, which corresponds to the critical size of the unstable droplet, obtained by setting $\rho_{v}=\rho$ and $\Phi=0$. When the system is confined (for instance $L=50$), the free energy shows two extrema: the first is a maximum associated with the unstable state and the second is a minimum that gives the size of the stable droplet. When the system is too small, the free energy of the second extrema becomes larger than that of the initial state (as for $L=25$) and nucleation is impeded due to superstabilization. Evidently, a prior estimation of the stable equilibrium, if existing, will help in the choice of the initial droplet for designing efficient seeded simulations.
@@ 4 Results
We first examine the influence of the initial droplet in NVT seeded simulations for the condensation of Lennard-Jones particles. A liquid seed of density $\bar{\rho}_{l}$ is introduced with an initial radius $R$ in a box of size $L=50$ with global density $\rho=0.025$ at $T=0.8$. Mass conservation determines the vapor density of the initial state. Various sizes for the liquid seed are used, ranging from $R=2$ to $R=9.7$. The NVT simulations run for $2000\tau$. Fig. 2 reports the size of the liquid droplet with time. When the seed is supercritical ($R>3$), the stable equilibrium is reached. It corresponds to a droplet of size $R\sim 8.5$ in equilibrium with the vapor phase. This stable configuration, which refers to a minimum in free energy, is in very good agreement with CNT predictions, see Fig.1 for $L=50$. The smallest seed ($R=2$) is found to redissolve, which also aligns with CNT predictions (Fig.1). NVT runs suggest that the critical (unstable) radius is indeed comprised between $R=2$ and $R=3$ (Fig.2). Note that, after disappearance of the smallest seed ($R=2$), thermal fluctuations produce the stochastic emergence of small clusters made of roughly 101510-15 particles, detected in Fig.2. None of them are found to reach the critical size and enter the growth regime within the time window simulated in our NVT runs, as expected since supersaturation is low. These NVT seeded simulations demonstrate that the stable droplet is independent of the size of the initial seed, as long as it is chosen within the supercritical region ($R>3$ in this case). Nevertheless, the time for reaching this stable equilibrium is naturally size dependent, highlighting the benefit from initializing with a size close to the targeted equilibrium. In the following, CNT predictions from Eqs. (4), (5), and (6) will be employed to initialize the seeding simulations.
Figure 2: Temporal evolution of the liquid droplet radius resulting from NVT seeded simulations in a box of size $L=50$ with global density $\rho=0.025$ at $T=0.8$. Various sizes for the liquid seed are employed to initialize the NVT simulations, ranging from $R=2$ to $R=9.7$. All seed sizes reach that of the stable droplet (near $R\sim 8.5$), except the smallest seed that redissolves.
We then investigate the condensation of Lennard-Jones particles over the temperature range from $0.7$ to $1.1$. The two solutions of Eqs. (4), (5), and (6) are obtained using chemical potential computed from the EOS developed by Johnson, Zollweg, and Gubbins (JZG) [52]. The EOS are combined with a mean-field correction for the cut potential [52], using the same cutoff ($6.78$) as in the simulations. Within this temperature range, the phase diagram predicted by the JZG EOS is found to be in excellent agreement with that obtained from MD simulations of the two-phase system at equilibrium (not shown). The surface tension of the planar vapor/liquid interface at equilibrium is classically calculated from MD simulations using the Kirkwood pressure tensor method [61]. The results follow the expected power law [58], $\gamma=\gamma_{0}\left(1-T/T_{c}\right)^{a}$, with $\gamma_{0}=2.838$, $a=1.252$, and $T_{c}=1.305$ the critical temperature.
@@
Refer to caption(a)Refer to caption(b)Refer to caption(c)Refer to caption(d)Refer to caption(e)Refer to caption
@@
The smaller droplet (dashed lines) corresponds to the unstable equilibrium, while the larger droplet (solid lines) represents the stable cluster. As expected [48, 50, 51], the two branches merge for the density at which the superstabilization effect operates, i.e., no cluster forms since the vapor phase at this given box density becomes stable. For supersaturation between the binodal limit (vertical black dotted line) and this box density, nucleation is therefore impeded [48]. Predictions obtained from the ideal gas approximation are also shown: stable and unstable branches are represented by dash-dotted and dotted lines, respectively. They agree very well with the JZG EOS results, particularly for the upper branches corresponding to the stable clusters.
@@
In order to measure the cluster properties, we use Ovito [62] and first segment the liquid droplet by clustering analysis. A clustering distance, $d$, is defined, and particles forming a connected set — each within a distance $d$ of at least one other — are identified; the number of such particles are denoted $N_{l}^{*}$. The clustering analysis is not very sensitive to $d$ within the conventional range [1.2, 1.5][1.2,\,1.5] for Lennard-Jones droplets [8]; in our analysis we adopt $d=1.25$. Several methods available in Ovito’s MD standard analysis library can then be used to determine the cluster radius: (i) fitting the radial density profile, or its integral, in simulation to a modified sigmoidal function [9, 63], (ii) the convex hull method [62], or (iii) the radius of gyration [62]. The radius may also be estimated from the droplet volume, $R^{*}=(3N_{l}^{*}/(4\mathrm{\pi}\bar{\rho}_{l}))^{1/3}$, assuming the droplet is homogeneous with bulk liquid density $\bar{\rho}_{l}$ from the phase diagram. In our case, all methods yield very similar results, since the clusters are well defined, nearly spherical, and relatively large. For sake of clarity, only the results obtained from the droplet volume estimation are shown. The corresponding results are also shown in Fig. 3(a) as light-colored circles with error bars. The error bars, representing standard deviations in $R^{*}$, are too small to be visible. With three configurations ($L=30$, $50$, and $70$), stable clusters with radius ranging from $5$ to $16$ are obtained. Remarkably, the seeding results are reproduced very well by CNT predictions using either the JZG EOS or the ideal gas approximation. Fig. 3(b)-(e) present seeded simulation results at higher temperatures: $T=0.8$, $T=0.9$, $T=1$, and $T=1.1$. Agreement with CNT predictions using the JZG EOS remains excellent across this range. In contrast, the ideal gas approximation becomes unreliable at high temperature. Nevertheless, it provides sufficiently accurate estimates to serve as practical initial conditions for seeded simulations, owing to the simplicity of the corresponding CNT calculations.
@@
Refer to caption(a)Refer to caption(b)Refer to caption(c)Refer to caption(d)Refer to caption(e)
@@ 5 Summary
The seeding technique has been applied to investigate the condensation of Lennard-Jones particles and, at the same time, to test both CNT and the accuracy of widely used thermodynamic models for Lennard-Jones fluid. The properties of the critical clusters predicted by CNT are found to agree well with seeded simulation results when JZG EOS is employed. The KN EOS also performs relatively well. In contrast, caution is required when using DFT or the ideal gas approximation, as both exhibit significant inaccuracies at high temperatures ($T\geq 0.9$). Nevertheless, the ideal gas approximation in confined systems, as expressed in Eqs. (6) and (8), remains of very practical value. It provides a simple and effective framework for initializing and guiding NVT seeded simulations, especially when no thermodynamic model for chemical potential is available. We restrict ourselves to the gas-liquid phase transition but the seeding technique was proved to be also a powerful technique for studying crystal nucleation [31, 35]. Coupled with CNT calculations, NVT seeded simulations are anticipated to be of great interest for quantitative investigation of materials provided realistic interatomic potentials exist.